OpenStudy (anonymous):

Find the derivative of tan(x^3-3x)

OpenStudy (zehanz):

YOu have to use the Chain Rule. Set u=x³-3x, then calculate d(tan(u))/du * du/dx

OpenStudy (campbell_st):

you need to use the chain rule \[\frac{dy}{dx} = \frac{dy}{du} \times \frac{du}{dx}\]

OpenStudy (zehanz):

@maink2: two people saying the same thing. Must be true! ;)

OpenStudy (anonymous):

The answer I got was Sec^2(3x^2-3)d/dx(x^3-3x).

OpenStudy (zehanz):

So what is d/dx(x^3-3x)?

OpenStudy (anonymous):

(3x^2-3) :)

OpenStudy (zehanz):

Now you're talkin'!

OpenStudy (zehanz):

BTW, I would have got:\[\frac{ 3x^2-3 }{ \cos^2x }\]because we're not into sec's and csc's in Europe...